OpenStudy (anonymous):

36+x=15 plz help will give medal

OpenStudy (xxdragoonxx):

36+x=15 36+x-36=15-36 x=-21

TheSmartOne (thesmartone):

\(\Huge\cal\color{blue}{Welcome~to~Open~Study!!}\) It looks like you are new here. This tutorial can help you get started. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8